# From Single to Multi: How LLMs Hallucinate in Multi-Document Summarization

### v1

@@ 5 What type of mistakes do LLMs make?
Instruction Inconsistency Insights that are unrelated to the subtopic, redundant, or not shared, violating the conditions specified in the prompt: (1) focus on a subtopic, (2) create concise summaries, and, in the “subtopic+shared” setting, (3) focus on shared information.
@@
Understanding the nature and frequency of LLM hallucinations is crucial for evaluating the reliability of these models in practical applications. While prior work acknowledges failures in LLM outputs (Laban et al., 2024), these failures are not well-studied in MDS. To bridge this gap, we collect human annotations for 150+ LLM-generated summaries and proposed a taxonomy based on the recurring mistakes observed. Specifically, we manually inspect over 700 predicted insights across models, domains, and task focus. To assess whether predicted insights are faithful, we also include the input documents in the annotation process. However, as previously observed (Chang et al., 2024), analyzing long documents is quite challenging. Consequently, instead of analyzing multiple combination sizes, we limit our analysis to N=2, leaving the analysis of 2+ documents for future work. The main analysis is conducted by two authors of this paper (see Appendix F for further details on the annotation protocol). Finally, to ensure the robustness of the proposed error categories, we collect the category annotations of two additional authors across 50 examples spanning both domains for each category (Chang et al., 2024). We discover that, on average, 80.71% of the annotated insights are assigned the same label by 3 or more annotators.
@@
Table 2 shows each error’s frequency in the news domain. Overall, we find all LLMs to be fairly faithful to input documents, as emphasized by the low rates of context inconsistency (9%-37%) and fabrication errors (0%-9%). The majority of the errors seems to be related to the generation of uninformative insights (pedantic) and unrelated subtopics (instruction inconsistency). For instance, GPT-4o and Qwen 2 (72B) often include redundant and off-topic insights, while also adding coherence-enhancing insights such as takeaway statements. In “subtopic+shared” results, the rate of instruction inconsistency errors exceeds 70% of all hallucinations. Of these, 80% to 95.45% (58.54% to 67.24% of all hallucinations) stem from insights not shared across documents, as required by the prompt.
Table 2: Observed hallucinations (and their frequency) when summarizing news articles for different task focus (“subtopic” and “subtopic+shared”). Reported values represent the fraction of error mistakes that are categorized as either Pedantic, Instruction Inconsistency, Context Inconsistency, and Fabrication. We also report the Total number of errors for each LLM.
@@ 6 Mitigating Hallucination
Across mitigation methods, F1-score variation is minimal ($\pm$3%), with a simple top-k showing the most improvement (see Table 3). Additionally, st-unrelated and redundant have less impact than expected, despite targeting common errors in generated summaries. This may be due to poor performance of adopted LLM-based classifiers in the tested domains. Even with a narrow focus (“subtopic”, N=2), our findings highlight the complexity of reducing hallucinated errors in LLM-generated summaries within a MDS setting, calling for further research to address these challenges.
Table 3: Absolute difference in average F1-score after applying four simple mitigation methods to summaries generated from two input documents (N=2). All methods show minimal impact on average F1-score ($\pm$3%), with truncating summaries to the top 5 bullet points being the most effective.
@@ 7 Related Work
Multi-document summarization is a broad and versatile multi-document NLP task that consists of generating a summary from multiple source documents, including opinion/reviews (Angelidis et al., 2021b; Iso et al., 2022), scientific articles (Lu et al., 2020; DeYoung et al., 2021; Yang et al., 2023), or news articles (Fabbri et al., 2019). Assessments of LLMs capabilities can be carried using generic summarization benchmarks (Fabbri et al. (2019); Lu et al. (2020); DeYoung et al. (2021); Wolhandler et al. (2022); Huang et al. (2024), inter alia) or using focus-specific benchmarks, that resort to aspects (Angelidis et al. (2021b); Hayashi et al. (2021); Amar et al. (2023); Yang et al. (2023), inter alia), queries (Kulkarni et al. (2020); Bolotova-Baranova et al. (2023); Huang et al. (2024); Liu et al. (2024b); Chen et al. (2024); Laban et al. (2024), inter alia), or perspectives (Naik et al., 2024). However, these benchmarks mainly focus on overall performance metrics and rarely conduct in-depth error analysis, offering limited insight into LLMs behavior and its relation to the input.
Hallucination evaluation benchmarks have been proposed for single-document NLP tasks with the purpose of evaluating models’ propensity to generate hallucinated content (Huang et al., 2023). These benchmarks are often designed to probe model errors, for instance, by testing knowledge of false beliefs and misconceptions (Lin et al., 2022; Cheng et al., 2023), current events (Vu et al., 2024; Kasai et al., 2024), similar but incorrect statements (Muhlgay et al., 2024), or even a models’ ability to recover from nonsensical questions (Pal et al., 2023). While single-document hallucination benchmarks have been crucial in identifying and mitigating non-factual outputs in current LLMs (Touvron et al., 2023; Yang et al., 2024; Achiam et al., 2024; Wei et al., 2022), hallucinatory behavior in multi-document tasks remains under-explored, especially how it varies with factors like document count or repeated or contradictory information.
@@ 8 Conclusion
In this paper, we carry the first in-depth investigation of hallucinatory behavior of popular instruction-tuned LLMs in a multi-document summarization task. Through controlled experiments where we vary the number of documents and task focus, we find that up to 75% of LLM-generated content is hallucinated. More surprisingly, our results show that when prompting models to summarize information related to non-existing subtopics, models still generate plausible-sounding summaries in more than 20% of the examples, with GPT-3.5-Turbo incorrectly generating summaries in 79.35% of the samples. Subsequent manual analysis of 700+ insights reveals that, albeit faithful to the input, a large fraction of insights is overly generic or inconsistent with the instructions. To mitigate such errors, we experiment with simple post-processing mitigation strategies but observe a trade-off with models’ performances. Our results underscore the need for more effective approaches that systematically mitigate hallucinations in MDS.
@@
While we design our evaluation protocol to ensure the reliability of our analysis, we acknowledge a few limitations with our work, which we list below:
@@
Insights may not represent a single unit of information. Underlying our evaluation protocol was the assumption that insights—both predicted and reference—represent a unique unit of information and “are expected to mention a number, a date, or an entity” and “are independent of each other” (Laban et al., 2024). However, we observe that the assumption does not always hold in practice (see examples in Appendix A), which has repercussions for the evaluation metric, which is produces a 1-to-1 coverage mapping between lists of insights. To address this, one solution could be to modify the evaluation prompt (listed in Figure 14) to output a list of “coverage” judgments (as opposed to a single JSON object). However doing so may introduce additional challenges for evaluation, such as anchoring effects for multi-attribute judgments (Stureborg et al., 2024). Alternatively, we could decompose the reference and predicted insights into smaller units of information (Min et al., 2023; Wanner et al., 2024) and apply our metric on such decomposed facts.
@@
The use of information coverage as a proxy for measuring similarity of meaning. We acknowledge the pragmatic and semantic nuances involved in determining whether two insights have the same meaning. Linguistic phenomena, such as entailment, implicature, and presupposition (Dahlman, 2021; Jeretic et al., 2020; Jiang and de Marneffe, 2021, 2019) could prove useful in determining the semantic relationships between two sentences (Goyal and Durrett, 2020). Alternatively, researchers have also explored other approaches (Tang et al., 2023a), based on iterative question generation and response (Wang et al., 2020; Scialom et al., 2021) or even based on prompting LLM using a few demonstrations (Chang et al., 2024; Laban et al., 2024). In this work, similarly to previous work (Huang et al., 2024; Laban et al., 2024), we use the more abstract notion of coverage (as opposed to other linguistic phenomena) due to its previous reports of high agreement with both experts and crowdsourcing annotators in the context of abstractive summarization tasks.
@@
Hallucination Mitigation Trade-off. In this paper, we explore the efficacy of simple heuristics in the mitigation of hallucinations across LLMs, which were designed to tackle specific issues with the LLM-generated summaries, including redundancy (Xiao and Carenini, 2020). However, we observe limited effect in LLMs’ hallucination rate in practice, raising important questions about the efficacy of these methods. One potential explanation is that the LLM-based classifiers, used to detect these errors, perform poorly in the specific domains tested in this work. Moreover, we observe that most insights (both reference and predicted) are rarely atomic, containing a mixture of redundant and non-redundant pieces of information, making it difficult to detect using binary judgements. Future work could explore the use of more fine-grained atomic evaluation methodologies (Min et al., 2023; Chen et al., 2023; Tang et al., 2024).
@@ Acknowledgments
The authors would like to thank the members of Megagon Labs, and the UCI-NLP group at UC Irvine for the provided feedback and insightful discussions regarding this project.
@@ Appendix B Prompt Selection
  • Figure 12, used in the news, “subtopic+shared” setting;
@@
  • Figure 13, used in the conv, “subtopic+shared” setting;
@@
Early on in our experiments, we explored a modified version of the news summarization prompt proposed in the SummHay paper (Laban et al., 2024). Since our goal was to investigate hallucinations in MDS (and not in model’s capability to cite different sources), we remove the instructions related to the citation task, therefore avoiding potential anchoring effects that could be associated with performing multiple tasks in a single shot (Stureborg et al., 2024). Additionally, because we are interested in analyzing model behavior in scenarios where the number of relevant insights in the input documents is unknown, we do not instruct the model on the ideal summary length. Instead, in our earlier experiments, to avoid biasing evaluation based on the length of the summaries (Liu et al., 2023; Guo and Vosoughi, 2023), we explore instructing the model to restrict the number of words used in the summaries. To determine this number, we analyse the ground truth summaries (concatenation of reference insights for each dataset) and find 300 words to be a reasonable upper bound over the reference insights (see Figure 7), therefore allowing for some semantic and lexical variations in LLM summaries.
Figure 7: Distribution of the number of words of the reference summaries for the proposed evaluation benchmark in the news domain. The length of reference summaries is computed by concatenating ground truth reference insights for the different examples in the evaluation benchmarks and counting the number of words separated by whitespaces. Any, and Shared refer to the “subtopic” and “subtopic+shared” settings, respectively. We observe that, across the different document combinations, the number of words (including punctuation) is fewer than the limit of 300 words specified in the prompt proposed by Laban et al. (2024). Thus suggesting that 300 words is more than enough to faithfully summarize information within the provided documents.
@@
Figure 8: Number of words in summaries as a function of the number of input documents (news domain). On the left, we observe the length of LLM-generated summaries when instructed to summarize insights in the “subtopic” setting. On the right, we observe the length of LLM-generated summaries when instructed to summarize shared subtopic-related insights (“subtopic+shared” setting). In general, we observe a large gap between the total length of ground truth summaries (black solid line) and the LLM-generated ones, suggesting that models attempt to generate 300 words-long summaries even when the relevant information can be condensed in using fewer words.
@@
Finally, to provide some format to the summaries and limit potential confounding aspects related to the bullet-point or the inability to identify subtopic-related insights in the prompt, we include two additional instructions in the prompt: Represent each bullet point using "-". and If you do not find related insights, write "No insights found" and nothing else..
@@
News Summarization - “subtopic+shared” Prompt You are given {{n_articles}} news articles about the main subject "{{topic}}". ``` Article 1: {{article_1}} Article 2: {{article_2}} {{remaining_articles}} ``` Your objective is to clearly and concisely summarize all insights regarding the topic of "{{subtopic}}" that are mentioned in at least two articles. In other words, your summary should include any insight that is related to the topic and is  mentioned in two or more articles. If you don’t find insights regarding the topic of "{{subtopic}}", return "No insights found". Careful: - [Format] You should format your summary as a bullet point list, where each bullet point is a different insight consisting of a single sentence. Represent each bullet point using "-". If you don’t find related insights, write "No insights found" and nothing else. - [Length] Your summary should be concise and clear.
Figure 12: Prompt used to summarize commonalities across news articles. The “subtopic+shared” prompt instructs the model to summarize insights that are shared across two or more documents and that relate to the specified {{subtopic}}. {{placeholders}} will be replaced with corresponding values.
Conv Summarization - “subtopic+shared” Prompt You are given {{n_articles}} conversations about the following scenario: "{{topic}}". The conversations involve the following participants: {{participants}}. In each conversation, the participants might be different, with different names, but all the conversations fall into the same scenario. ``` Conversation 1: {{conversation_1}} Conversation 2: {{conversation_2}} {{remaining_conversations}} ``` Your objective is to clearly and concisely summarize all insights regarding the topic of "{{subtopic}}" that are mentioned in at least two conversations. In other words, your summary should include any insight that is related to the topic and is mentioned in two or more conversations. If you don’t find insights regarding the topic of "{{subtopic}}", write "No insights found". Careful: - [Format] You should format your summary as a bullet point list, where each bullet point is a different insight consisting of a single sentence. Represent each bullet point using "-". If you don’t find related insights, write "No insights found" and nothing else. - [Length] Your summary should be concise and clear.
Figure 13: Prompt used to summarize commonalities across conversations. The “subtopic+shared” prompt instructs the model to summarize insights that are shared across two or more documents and that relate to the specified {{subtopic}}. {{placeholders}} will be replaced with corresponding values.

### v2

@@ 5 What type of mistakes do LLMs make?
Instruction Inconsistency Insights that are unrelated to the subtopic, redundant, or not shared, violating the conditions specified in the prompt: (1) focus on a subtopic, (2) create concise summaries, and, in the “subtopic+trustworthy” setting, (3) focus on shared information.
@@
Understanding the nature and frequency of LLM hallucinations is crucial for evaluating the reliability of these models in practical applications. While prior work acknowledges failures in LLM outputs (Laban et al., 2024), these failures are not well-studied in MDS. To bridge this gap, we collect human annotations for 150+ LLM-generated summaries and proposed a taxonomy based on the recurring mistakes observed. Specifically, we manually inspect over 700 predicted insights across models, domains, and task focus. To assess whether predicted insights are faithful, we also include the input documents in the annotation process. However, as previously observed (Chang et al., 2023), analyzing long documents is quite challenging. Consequently, instead of analyzing multiple combination sizes, we limit our analysis to N=2, leaving the analysis of 2+ documents for future work. The main analysis is conducted by two authors of this paper (see Appendix F for further details on the annotation protocol). Finally, to ensure the robustness of the proposed error categories, we collect the category annotations of two additional authors across 50 examples spanning both domains for each category (Chang et al., 2023). We discover that, on average, 80.71% of the annotated insights are assigned the same label by 3 or more annotators.
@@
Table 2 shows each error’s frequency in the news domain. Overall, we find all LLMs to be fairly faithful to input documents, as emphasized by the low rates of context inconsistency (9%-37%) and fabrication errors (0%-9%). The majority of the errors seems to be related to the generation of uninformative insights (pedantic) and unrelated subtopics (instruction inconsistency). For instance, GPT-4o and Qwen 2 (72B) often include redundant and off-topic insights, while also adding coherence-enhancing insights such as takeaway statements. In “subtopic+trustworthy” results, the rate of instruction inconsistency errors exceeds 70% of all hallucinations. Of these, 80% to 95.45% (58.54% to 67.24% of all hallucinations) stem from insights not shared across documents, as required by the prompt.
Table 2: Observed hallucinations (and their frequency) when summarizing news articles for different task focus (“subtopic” and “subtopic+trustworthy”). Reported values represent the fraction of error mistakes that are categorized as either Pedantic, Instruction Inconsistency, Context Inconsistency, and Fabrication. We also report the Total number of errors for each LLM.
@@ 6 Mitigating Hallucination
Across mitigation methods, F1-score variation is minimal ($\pm$3%), with a simple top-k showing the most improvement (see Table 3). Additionally, st-unrelated and redundant have less impact than expected, despite targeting common errors in generated summaries. This may be due to poor performance of adopted LLM-based classifiers in the tested domains. Even with a narrow focus (“subtopic”, N=2), our findings highlight the complexity of reducing hallucinated errors in LLM-generated summaries within a MDS setting, calling for further research to address these challenges. To minimize hallucination in MDS, future work could explore hierarchical approaches, where each document is summarized individually before combining the results into a final summary (Chang et al., 2023; Tang et al., 2024).
Table 3: Absolute difference in average F1-score after applying four simple mitigation methods to summaries generated from two input documents (N=2). All methods show minimal impact on average F1-score ($\pm$3%), with truncating summaries to the top 5 bullet-points being the most effective.
@@ 7 Related Work
Multi-document summarization is a broad and versatile multi-document NLP task that consists of generating a summary from multiple source documents, including opinion/reviews (Angelidis et al., 2021b; Iso et al., 2022), scientific articles (Lu et al., 2020; DeYoung et al., 2021; Yang et al., 2023), or news articles (Fabbri et al., 2019). Assessments of LLMs capabilities can be carried using generic summarization benchmarks (Fabbri et al. (2019); Lu et al. (2020); DeYoung et al. (2021); Wolhandler et al. (2022); Huang et al. (2024), inter alia) or using focus-specific benchmarks, that resort to aspects (Angelidis et al. (2021b); Hayashi et al. (2021); Amar et al. (2023); Yang et al. (2023), inter alia), queries (Kulkarni et al. (2020); Bolotova-Baranova et al. (2023); Huang et al. (2024); Chen et al. (2024); Laban et al. (2024), inter alia), or perspectives (Naik et al., 2024). However, these benchmarks mainly focus on overall performance metrics and rarely conduct in-depth error analysis, offering limited insight into LLMs behavior and its relation to the input.
Hallucination evaluation benchmarks have been proposed for single-document NLP tasks with the purpose of evaluating models’ propensity to generate hallucinated content (Huang et al., 2025). These benchmarks are often designed to probe model errors, for instance, by testing knowledge of false beliefs and misconceptions (Lin et al., 2022; Cheng et al., 2023), current events (Vu et al., 2024; Kasai et al., 2023), similar but incorrect statements (Muhlgay et al., 2024), or even a models’ ability to recover from nonsensical questions (Pal et al., 2023). While single-document hallucination benchmarks have been crucial in identifying and mitigating non-factual outputs in current LLMs (Touvron et al., 2023; Yang et al., 2024; Achiam et al., 2024; Wei et al., 2022), hallucinatory behavior in multi-document tasks remains under-explored, especially how it varies with factors like document count or repeated or contradictory information.
Automatic evaluation, while less reliable than human evaluation (Kryscinski et al., 2019; Fabbri et al., 2021), is often sought for its scalability and efficiency in large-scale experiments, where human evaluation is impractical (Krishna et al., 2023; Chang et al., 2023; Huang et al., 2024). To evaluate the grounding of LLM-generated text, specialized entailment-based (Laban et al., 2022; Tang et al., 2024; Goyal and Durrett, 2021) and question-answering-based (Fabbri et al., 2022) metrics have been proposed. Recent work has found that, despite their higher cost, zero-shot prompting general-purpose LLMs performs as well as or better than specialized metrics when used for fact-checking while requiring no fine-tuning(Manakul et al., 2023; Tang et al., 2024).666 Similarly, prompting LLMs (with task specific prompts) has been shown to correlate better with human judgments than specialized models (Liu et al., 2023a; Farquhar et al., 2024). Building on prior successes of LLM-as-a-judge approaches, we use an LLM-based metric with three demonstration examples (Chang et al., 2023; Laban et al., 2024) to determine the information coverage between reference and predicted insights. While a variant of this metric has been rigorously validated (Laban et al., 2024), we further confirm its reliability through human evaluations (see Sections 3 and 5), finding strong alignment with LLM judgments.
@@ 8 Conclusion
In this paper, we carry the first in-depth investigation of hallucinatory behavior of popular instruction-tuned LLMs in a multi-document summarization task. Through controlled experiments where we vary the number of documents and task focus, we find that up to 75% of LLM-generated content is hallucinated. More surprisingly, our results show that when prompting models to summarize information related to non-existing subtopics, models still generate plausible-sounding summaries in more than 20% of the examples, with GPT-3.5-Turbo incorrectly generating summaries in 79.35% of the samples. Subsequent manual analysis of 700+ insights reveals that, albeit faithful to the input, a large fraction of insights is overly generic or inconsistent with the instructions. To mitigate such errors, we experiment with simple post-processing mitigation strategies but observe a trade-off with models’ performances. Together, our results underscore the need for more effective approaches that systematically mitigate hallucinations in MDS.
@@
While we design our evaluation protocol to ensure the reliability of our analysis, we acknowledge some limitations with our work.
Grounding evaluation in reference insights. To automate evaluation while keeping costs manageable, we compare LLM-predicted insights with the reference insights instead of using the documents. This assumes that the reference insights cover all relevant document content. This assumption is supported by the fact that SummHay documents were artificially generated from the reference insights and rigorously validated for completeness (Laban et al., 2024). However, this simplification could result in misclassifying a predicted insight as hallucinated, even if it is covered in the document. Empirically, across 700+ insights, we find that most “false hallucination” results from models merging two high-level ideas, rather than omitting important information. Alternatively, future work could explore document-level (Yin et al., 2021), sentence-level (Huang et al., 2025), or fact-level (Tang et al., 2024) approaches to directly cross-check predicted insights with the documents themselves, rather than with reference insights.
@@
Insights may not represent a single unit of information. Underlying our evaluation protocol was the assumption that insights—both predicted and reference—represent a unique unit of information and “are expected to mention a number, a date, or an entity” and “are independent of each other” (Laban et al., 2024). However, we observe that the assumption does not always hold in practice (see examples in Appendix A), which has repercussions for the evaluation metric, which is produces a 1-to-1 coverage mapping between lists of insights. To address this, one solution could be to modify the evaluation prompt (listed in Figure 15) to output a list of “coverage” judgments (as opposed to a single JSON object). However doing so may introduce additional challenges for evaluation, such as anchoring effects for multi-attribute judgments (Stureborg et al., 2024). Alternatively, we could decompose the reference and predicted insights into smaller units of information (Min et al., 2023; Wanner et al., 2024) and apply the metric on decomposed facts.
@@
The use of information coverage as a proxy for measuring similarity of meaning. We acknowledge the pragmatic and semantic nuances involved in determining whether two insights have the same meaning. Linguistic phenomena, such as entailment, implicature, and presupposition (Dahlman, 2021; Jeretic et al., 2020; Jiang and de Marneffe, 2021, 2019) could prove useful in determining the semantic relationships between two sentences (Goyal and Durrett, 2020). Alternatively, researchers have also explored other approaches (Tang et al., 2023a), based on iterative question generation and response (Wang et al., 2020; Scialom et al., 2021) or even based on prompting LLM using a few demonstrations (Chang et al., 2023; Laban et al., 2024). In this work, similarly to previous work (Huang et al., 2024; Laban et al., 2024), we use the more abstract notion of coverage (as opposed to other linguistic phenomena) due to its previous reports of high agreement with both experts and crowdsourcing annotators for abstractive summarization tasks.
@@
Hallucination Mitigation Trade-off. In this paper, we explore the efficacy of simple heuristics in the mitigation of hallucinations across LLMs, which were designed to tackle specific issues with the LLM-generated summaries, including redundancy (Xiao and Carenini, 2020). However, we observe limited effect in LLMs’ hallucination rate in practice, raising important questions about the efficacy of these methods. One potential explanation is that the LLM-based classifiers, used to detect these errors, perform poorly in the specific domains tested in this work. Moreover, we observe that most insights (both reference and predicted) are rarely atomic, containing a mixture of redundant and non-redundant pieces of information, making it difficult to detect using binary judgments. Future work could explore the use of more fine-grained atomic evaluation methodologies (Min et al., 2023; Chen et al., 2023b; Tang et al., 2024).
@@ Acknowledgments
We thank the members of Megagon Labs and UCI NLP for their valuable feedback and fruitful discussions.
@@ Appendix B Prompt Selection
  • Figure 13, used in the news, “subtopic+trustworthy” setting;
@@
  • Figure 14, used in the conv, “subtopic+trustworthy” setting;
@@
Early on in our experiments, we explored a modified version of the news summarization prompt proposed in the SummHay paper (Laban et al., 2024). Since our goal was to investigate hallucinations in MDS (and not in model’s capability to cite different sources), we remove the instructions related to the citation task, therefore avoiding potential anchoring effects that could be associated with performing multiple tasks in a single shot (Stureborg et al., 2024). Additionally, because we are interested in analyzing model behavior in scenarios where the number of relevant insights in the input documents is unknown, we do not instruct the model on the ideal summary length. Instead, in our earlier experiments, to avoid biasing evaluation based on the length of the summaries (Liu et al., 2023b; Guo and Vosoughi, 2023), we explore instructing the model to restrict the number of words used in the summaries. To determine this number, we analyse the ground truth summaries (concatenation of reference insights for each dataset) and find 300 words to be a reasonable upper bound over the reference insights (see Figure 8), therefore allowing for some semantic and lexical variations in LLM summaries.
Figure 8: Distribution of the number of words of the reference summaries for the proposed evaluation benchmark in the news domain. The length of reference summaries is computed by concatenating ground truth reference insights for the different examples in the evaluation benchmarks and counting the number of words separated by whitespaces. Any, and Shared refer to the “subtopic” and “subtopic+trustworthy” settings, respectively. We observe that, across the different document combinations, the number of words (including punctuation) is fewer than the limit of 300 words specified in the prompt proposed by Laban et al. (2024). Thus suggesting that 300 words is more than enough to faithfully summarize information within the provided documents.
@@
Figure 9: Number of words in summaries as a function of the number of input documents (news domain). On the left, we observe the length of LLM-generated summaries when instructed to summarize insights in the “subtopic” setting. On the right, we observe the length of LLM-generated summaries when instructed to summarize shared subtopic-related insights (“subtopic+trustworthy” setting). In general, we observe a large gap between the total length of ground truth summaries (black solid line) and the LLM-generated ones, suggesting that models attempt to generate 300 words-long summaries even when the relevant information can be condensed in using fewer words.
@@
Finally, to provide some format to the summaries and limit potential confounding aspects related to the bullet-point or the inability to identify subtopic-related insights in the prompt, we include two additional instructions in the prompt: Represent each bullet-point using "-". and If you do not find related insights, write "No insights found" and nothing else..
@@
News Summarization - “subtopic+trustworthy” Prompt You are given {{n_articles}} news articles about the main subject "{{topic}}". ``` Article 1: {{article_1}} Article 2: {{article_2}} {{remaining_articles}} ``` Your objective is to clearly and concisely summarize all insights regarding the topic of "{{subtopic}}" that are mentioned in at least two articles. In other words, your summary should include any insight that is related to the topic and is  mentioned in two or more articles. If you don’t find insights regarding the topic of "{{subtopic}}", return "No insights found". Careful: - [Format] You should format your summary as a bullet point list, where each bullet point is a different insight consisting of a single sentence. Represent each bullet point using "-". If you don’t find related insights, write "No insights found" and nothing else. - [Length] Your summary should be concise and clear.
Figure 13: Prompt used to summarize commonalities across news articles. The “subtopic+trustworthy” prompt instructs the model to summarize insights that are shared across two or more documents and that relate to the specified {{subtopic}}. {{placeholders}} will be replaced with corresponding values.
Conv Summarization - “subtopic+trustworthy” Prompt You are given {{n_articles}} conversations about the following scenario: "{{topic}}". The conversations involve the following participants: {{participants}}. In each conversation, the participants might be different, with different names, but all the conversations fall into the same scenario. ``` Conversation 1: {{conversation_1}} Conversation 2: {{conversation_2}} {{remaining_conversations}} ``` Your objective is to clearly and concisely summarize all insights regarding the topic of "{{subtopic}}" that are mentioned in at least two conversations. In other words, your summary should include any insight that is related to the topic and is mentioned in two or more conversations. If you don’t find insights regarding the topic of "{{subtopic}}", write "No insights found". Careful: - [Format] You should format your summary as a bullet point list, where each bullet point is a different insight consisting of a single sentence. Represent each bullet point using "-". If you don’t find related insights, write "No insights found" and nothing else. - [Length] Your summary should be concise and clear.
Figure 14: Prompt used to summarize commonalities across conversations. The “subtopic+trustworthy” prompt instructs the model to summarize insights that are shared across two or more documents and that relate to the specified {{subtopic}}. {{placeholders}} will be replaced with corresponding values.